# Seq-VCR:Preventing Collapse in Intermediate Transformer Representations for Enhanced Reasoning

### v1

@@ Abstract
Decoder-only Transformers often struggle with complex reasoning tasks, particularly arithmetic reasoning requiring multiple sequential operations. In this work, we identify representation collapse in the model’s intermediate layers as a key factor limiting their reasoning capabilities. To address this, we propose Sequential Variance-Covariance Regularization (Seq-VCR [^1] https://github.com/rarefin/seq_vcr), which enhances the entropy of intermediate representations and prevents collapse. Combined with dummy pause tokens as substitutes for chain-of-thought (CoT) tokens, our method significantly improves performance in arithmetic reasoning problems. In the challenging $5\times 5$ integer multiplication task, our approach achieves $99.5\%$ exact match accuracy, outperforming models of the same size (which yield $0\%$ accuracy) and GPT-4 with five-shot CoT prompting ($44\%$). We also demonstrate superior results on arithmetic expression and longest increasing subsequence (LIS) datasets. Our findings highlight the importance of preventing intermediate layer representation collapse to enhance the reasoning capabilities of Transformers and show that Seq-VCR offers an effective solution without requiring explicit CoT supervision.
@@ 2 Literature Review
Representation collapse in representation learning degrades model performance by making features indistinguishable. It often occurs in unsupervised and self-supervised learning tasks due to the lack of labeled data, leading to trivial solutions and a low-rank feature space. Research by Jing et al. (2021) highlights that representation collapse is due to poor optimization, loss function design, and improper regularization. Recent studies (Grill et al., 2020) suggest the use of contrastive learning, such as SimCLR, to mitigate this by contrasting positive and negative samples. Additionally, Zbontar et al. (2021) introduced Barlow Twins to minimize redundancy between learned representations. A further advancement is VICReg (Bardes et al., 2021; Zhu et al., 2023) which employs variance, invariance, and covariance regularization to prevent collapse, ensuring diverse and non-trivial representations. While these researches focuse on images, it is less explored in language models. Recently, Barbero et al. (2024) studied last-layer representations of the next tokens and found representation collapse in pre-trained models.
@@ 3.3 Representation Collapse in Transformers
Figure 2: Representation collapse across layers during (a) training (or fine-tuning(b)) for two datasets. The x-axis represents the layer ID, while the y-axis shows the degree of collapse as measured by representation Matrix-Entropy. The results highlight how intermediate layers (shown by the decline in Entropy) experience representation collapse during Vanilla training or fine-tuning for both datasets, indicating potential bottlenecks in information flow or feature learning. Tools like Pause Goyal et al. (2023b) token-based tuning can’t fix it, but our proposed regularization Seq-VCR can improve collapse.
Representation collapse in Transformers manifests as a reduction in the diversity of internal representations across layers, particularly in tasks requiring complex reasoning. This collapse hinders the model’s ability to capture and process intricate patterns necessary for tasks like multi-digit multiplication.
To illustrate this, we analyze the layer-wise representation entropy of a GPT-2 model fine-tuned on the $5\times 5$ digit multiplication task. As shown in Figure 2(b), there is a significant drop in entropy in the intermediate layers, indicating a collapse in representation diversity. We hypothesize that this representation collapse is the reason why it is very hard for the model to calculate multiplication, carry, and store them, thus resulting in poorer performance, as shown in Figure 6 for the output positions, which require more computation (Figure 1).
For computing the representation collapse of a token sequence, we utilise the matrix entropy of the representation vectors in different layers of the transformer model. We investigate $\alpha$-order matrix-based entropy (Giraldo et al., 2014; Skean et al., 2023) as a practical alternative to Rényi’s $\alpha$-order entropy (Rényi, 1961). This metric uses a similarity kernel $\kappa$ on layer-wise representations of token sequences from a distribution, without assuming the true nature of the distribution. The kernel selection $\kappa$ is flexible, including Gaussian, linear, or Laplacian kernels. We choose the linear kernel $\kappa(a,b)=ab^{T}$ due to the linear representation hypothesis (Park et al., 2024), suggesting that language model representations encode high-level concepts in linearly separable manifolds.
Aussuming $Z^{(l)}=f_{l}\in\mathbb{R}^{T\times d}$ is the $l^{th}$ representations of the $T$ tokens in the sequence of dimension $d$, we can first construct the Gram matrix $\mathbf{K}=\kappa\left(Z^{(l)},Z^{(l)}\right)=Z^{(l)}Z^{(l)^{T}}\in\mathbb{R}^{\textrm{T}\times\textrm{T}}$ consisting of all pairwise evaluations of the $T$ token representations in $Z^{(l)}$. Then the matrix-based entropy of order $\alpha>0$ is defined as:
@@
Equation 2 is the $\alpha$-order Rényi entropy of the Gram matrix eigenvalues [^2] The non-zero eigenvalues of the Gram matrix $Z^{(l)}Z^{(l)^{T}}$ are equivalent to those of the covariance matrix $Z^{(l)^{T}}Z^{(l)}$. Using the covariance matrix instead of the Gram matrix in Equation 2 makes no difference and is more computationally efficient if $d<T$.. Each eigenvalue is divided by $\textrm{tr}(\mathbf{K})$ and raised to the $\alpha$ power so that the eigenvalues of $\mathbf{K}$ sum to $1$, as $\textrm{tr}(\cdot)=\sum_{i=1}^{n}\lambda_{i}(\cdot)$), allowing them to be treated as a probability distribution. Each eigenvalue of $\mathbf{K}$ represents the variance of the sequence token representations in a specific direction of the principal component Scholkopf & Smola (2018). Low entropy indicates a heavy-tailed distribution in which few components dominate the variance in $Z^{(l)}$. At maximum entropy, the eigenvalues form a uniform distribution, spreading token representations equally in all directions. Depending on $\alpha$, various special cases of matrix-based entropy can be recovered. Specifically, when $\lim_{\alpha\rightarrow 1}$ it equals Shannon entropy, also known as Von Neumann entropy in quantum information theory Bach (2022); Boes et al. (2019)).
@@ 3.4 Sequential Variance Covariance Regularization (Seq-VCR)
To prevent representation collapse, we propose Seq-VCR, a regularization technique that encourages the model to maintain high variance and low covariance in its intermediate representations. Seq-VCR is inspired by techniques used in self-supervised learning for image representations (Bardes et al., 2021) and is adapted here for sequential data in Transformer models.
Seq-VCR is applied to the final output of the model ($X=f_{cls}$). Since it requires a calculation of the covariance matrix and a pre-trained model such as GPT-2, the dimension can be $f_{cls}\mathbb{R}^{5000\times d}$, such dimensional calculation is often difficult. So we use a linear projection layer of the representation layer $f_{proj}$ which projects the representation of the layer $l$ into a smaller embedding space such as $X=f_{proj}(f_{l})R^{T\times d}$, $f_{proj}\in\mathbb{R}^{2048\times d}$.
Now we can assume that we have $N$ samples in the batch, so $N$ number of $X$. We can calculate the covariance matrix across the batch dimension, $C\in R^{T\times d\times d}$.
@@
where $\lambda_{1}$ and $\lambda_{2}$ are the coefficients of the variance and covariance regularization terms. $\eta$ is a small constant, used for numerical stability.
The Variance Term encourages unit variance in each dimension, while the Covariance Term penalizes covariance between different dimensions, promoting decorrelation and diversity in representations.
@@ 3.5 Incorporating Pause Tokens
We further enhance the model’s reasoning capacity by introducing dummy pause tokens, which act as placeholders for intermediate computation steps. Unlike explicit CoT tokens, pause tokens do not require supervision and serve to allocate additional computational resources within the model (Goyal et al., 2023b).
@@ 3.6 Training Objective
This encourages the model to not only predict the next token accurately but also maintain diverse and informative intermediate representations.
@@ 4.1 Data
Finally, we also conduct experiments in a more general setting outside arithmetic tasks, called Dynamic Programming (DP). Dynamic Programming is a framework for solving decision-making problems. The core idea in this setting is to break down a complex task into a series of smaller subtasks that can be solved sequentially. We chose a very popular problem in this setting called the Longest Increasing Sub-sequence (LIS) as described in the Introduction to Algorithms book (Cormen et al., 2022). For this task, the goal is to find the length of the longest increasing subsequence of a given integer sequence. We generate datasets with different input sequence lengths ranging from {50, 80, 100} as shown in Feng et al. (2024). Moreover, all input sequences, and answers in LIS are bounded-range integers and can therefore be tokenized (similar to the multi-digit multiplication task). The CoT tokens for this task consist of the dynamic programming array plus the final answer. Figure 3 shows a working example for a sample problem in all the datasets considered.
Moreover, Table 3 in the Appendix provides an overview of different tasks across the datasets i.e., Multiplication, Arithmetic Expression and LIS, along with their respective input, chain-of-thought (CoT), and output token details.
@@ 4.2 Models and Training Configurations
  • Vanilla: Standard training without CoT or pause tokens.
  • With CoT: Training with explicit CoT supervision.
@@ 4.3.1 Representation Diversity
Seq-VCR significantly enhances representation diversity, as evidenced by the layer-wise distribution of representation entropy across different configurations. In Figure 4, we observe that configurations employing our regularization technique, particularly Seq-VCR and Seq-VCR+Pause, exhibit higher entropy values in intermediate layers compared to the other configurations (Vanilla, and Pause). This increased diversity in representation entropy suggests that the model is able to capture a wider range of features and maintain distinct representations. The histograms also reveal that while other configurations show peaks at lower entropy values, indicating potential representation collapse and redundancy in the intermediate representations, our method fosters a more uniform distribution across the layers. This uniformity suggests a richer feature space being utilized by the model, enabling it to generalize better to unseen data, thus improving performance on downstream tasks.
@@ 4.4 Learning Dynamics
Our regularization method, Seq-VCR or Seq-VCR+Pause cause a significant phase transition (improving the next token prediction loss) in the performance of the model across datasets compared to other configurations such as Vanilla, and Pause, which do not induce this phase transition as shown in Figure 5. Specifically, this phase transition marks a distinct shift in the model’s ability to capture and generalize representations during training or finetuning.
Without regularization, learning is difficult and training loss saturates. Seq-VCR improves the learning curve. Adding Seq-VCR with $2$ Pause tokens causes a sharp phase transition (Figure 5 (a)) and solves the $5\times 5$ digit multiplication task. Figure 5 (b) shows similar improvements in arithmetic reasoning task that is learned from scratch.
@@ 4.5 Results on $4\times 4$ and $5\times 5$ digits Multiplication Tasks
We also see from Table 1 that GPT-4 with CoT 5-shot prompts achieves $44.0\%$ for $5\times 5$ digit multiplication, while Vanilla, Seq-VCR, and Pause achieve $0\%$ without CoT. Because, all configurations struggle with predicting middle tokens in the output sequence (shown in Figure 6). This happens because those middle tokens require more computations to calculate successive carry and store them, since these tokens require more computation (as shown in Figure 1), thus making them difficult sub-task than predicting the peripheral tokens which requires less computes, thus easier to solve. However, Seq-VCR+Pause with $2$ Pause tokens solves the $5\times 5$ digit multiplication task, previously unsolved without CoT tokens. Figure 2(b) shows our regularization Seq-VCR increases intermediate layer entropy, allowing more exploration in the representation space. We hypothesize that along with better representation by Seq-VCR, $2$ Pause tokens increase the computational ability, allowing the model to carry the necessary multiplication values and solve the task.
Figure 6: Position-wise accuracy of Multiplication of 5x5 digits with different configuration of GPT-2 Small. We see models fail on the middle tokens they require more compute.
@@ 4.6 Results on Arithmetic Expression and Longest Increasing Sub-sequence (LIS) Datasets
Figure 7 illustrates how the accuracy varies across different methods (represented by different bar colors in the plot) and with changing levels of task complexity, such as the number of operators in case of the Arithmetic Expressions Dataset or input sequence length in case of the LIS Dataset.
We can observe that as the task complexity is low i.e in case of 4 operators in the arithmetic expression and input sequence length of 50 in the LIS dataset, all the methods perform quite well. But, as the number of operators or sequence length increases, the tasks become more complex which shows the need for models other than the vanilla baselines.
Seq-VCR demonstrates substantial improvement over vanilla models and achieved results comparable to chain-of-thought (CoT) prompting method as shown in Feng et al. (2024). While CoT outperforms the other methods, there is an increase in the serial computation due to its step-by-step reasoning approach. On the other hand, Pause Tokens and regularization primarily boost parallel computation, enhancing the model’s expressivity for these types of problems without significantly increasing serial computation. The results presented in Figure 7 are based on models with a 5-layer configuration, optimized using the best-performing number of pause tokens. We find that the number of pause-tokens that works best for a particular model and layer config vary a bit. For example, the best working Seq-VCR + Pause model uses 7 pause tokens in case of the 5 operator task in Arithmetic Expression while 4 pause tokens works the best for the LIS task with 80 as input sequence length. Through Figure 8 we show how the different models perform by itself in the most complex task in each dataset. We also clearly see that more depth of the model helps every method in solving the task better. Another conclusion we can draw from figures 7 8 is that even though the combination of Seq-VCR with Pause Tokens works the best (almost close to CoT method), Seq-VCR gets similar test accuracy without the use of Pause Tokens in both datasets. All results in figure 78 are over 3 seed runs. Detailed ablation studies, exploring variations in the number of pause tokens and layers, can be found in the appendix.

### v2

@@ Abstract
Decoder-only Transformers often struggle with complex reasoning tasks, particularly arithmetic reasoning requiring multiple sequential operations. In this work, we identify representation collapse in the model’s intermediate layers as a key factor limiting their reasoning capabilities. To address this, we propose Sequential Variance-Covariance Regularization (Seq-VCR) [^1] https://github.com/rarefin/seq_vcr, which enhances the entropy of intermediate representations and prevents collapse. Combined with dummy pause tokens as substitutes for chain-of-thought (CoT) tokens, our method significantly improves performance in arithmetic reasoning problems. In the challenging $5\times 5$ integer multiplication task, our approach achieves $99.5\%$ exact match accuracy, outperforming models of the same size (which yield $0\%$ accuracy) and GPT-4 with five-shot CoT prompting ($44\%$). We also demonstrate superior results on arithmetic expression and longest increasing subsequence (LIS) datasets. Our findings highlight the importance of preventing intermediate layer representation collapse to enhance the reasoning capabilities of Transformers and show that Seq-VCR offers an effective solution without requiring explicit CoT supervision.
@@ 1 Introduction
The paper is structured as follows: Section 2 reviews reasoning in large language models, representation collapse, and regularization. Section 3 presents our method, Sequential Variance-Covariance Regularization (Seq-VCR), including its formulation and motivation. Section 4 describes the experimental setup, datasets, models, and shows Seq-VCR’s effectiveness in arithmetic reasoning and its impact on representations. Section 5 summarizes key findings and discusses future directions.
@@ 2 Literature Review
Representation collapse in representation learning degrades model performance by making features indistinguishable. It often occurs in unsupervised and self-supervised learning tasks due to the lack of labeled data, leading to trivial solutions and a low-rank feature space. Research by Jing et al. (2021) highlights that representation collapse is due to poor optimization, loss function design, and improper regularization. Recent studies (Grill et al., 2020) suggest the use of contrastive learning, such as SimCLR, to mitigate this by contrasting positive and negative samples. Additionally, Zbontar et al. (2021) introduced Barlow Twins to minimize redundancy between learned representations. A further advancement is VICReg (Bardes et al., 2021; Zhu et al., 2023) which employs variance, invariance, and covariance regularization to prevent collapse, ensuring diverse and non-trivial representations. While these researches focus on images, it is less explored in language models. Recently, Barbero et al. (2024) studied last-layer representations of the next tokens and found representation collapse in pre-trained models.
@@ 3.3 Representation Collapse in Transformers
Figure 2: Representation collapse across layers during (a) training (or fine-tuning(b)) for two datasets. The x-axis represents the layer ID, while the y-axis shows the degree of collapse as measured by representation Matrix-Entropy. The results highlight how intermediate layers (shown by the decline in Entropy) experience representation collapse for Pretrained GPT-2 Small on 5 × 5 digit Multiplication (b) and during Vanilla training or fine-tuning for both datasets, indicating potential bottlenecks in information flow or feature learning. Tools like Pause Goyal et al. (2023b) token-based tuning can’t fix it, but our proposed regularization Seq-VCR can improve collapse.
Representation collapse in Transformers manifests as a reduction in the diversity of internal representations across layers, particularly in tasks requiring complex reasoning. This collapse hinders the model’s ability to capture and process intricate patterns necessary for tasks like multi-digit multiplication. To illustrate this, we analyze the layer-wise representation entropy of a GPT-2 model fine-tuned on the $5\times 5$ digit multiplication task. As shown in Figure 2(b), there is a significant drop in entropy in the intermediate layers, indicating a collapse in representation diversity. We hypothesize that this representation collapse is the reason why it is very hard for the model to calculate multiplication, carry, and store them, thus resulting in poorer performance, as shown in Figure 6 for the output positions, which require more computation (Figure 1).
We quantify the phenomenon of representation collapse in token sequences by computing the matrix-based entropy of representation vectors across transformer model layers. We specifically explore the $\alpha$-order matrix-based entropy (Skean et al., 2023; Giraldo et al., 2014) using a similarity kernel $\kappa$ applied to layer-wise token representations without assuming the underlying distribution. We use a linear kernel defined as $\kappa(a,b)=ab^{T}$ in our analysis as in (Skean et al., 2024). Assuming, $Z^{(l)}=f_{l}\in\mathbb{R}^{T\times d}$ be the $l^{th}$ layer representations of $T$ tokens in a sequence with dimension $d$, we compute the pairwise similarity of $T$ token representations to form the Gram matrix $\mathbf{K}=\kappa\left(Z^{(l)},Z^{(l)}\right)=Z^{(l)}Z^{(l)^{T}}\in\mathbb{R}^{\textrm{T}\times\textrm{T}}$. Then, we determine the matrix-based entropy of order $\alpha>0$:
@@
[…]
@@ 3.4 Sequential Variance Covariance Regularization (Seq-VCR)
To address representation collapse, we introduce Seq-VCR, a regularization technique that enforces high variance and low covariance in intermediate representations of Transformer models. Adapted for sequential data, Seq-VCR builds directly on the variance and covariance regularization principles of VICReg (Bardes et al., 2021) and VCReg (Zhu et al., 2023), with Equation  3 re-purposing the formulations from VICReg to tackle the challenges of sequence modeling.
Seq-VCR is implemented on the final output of the model ($X=f_{cls}$) [^3] Initial tests indicated optimal performance when regularization was utilized on the final layer. Experimentation revealed that additional layers can also be effective. For detailed information on layer-specific regularization, refer to Appendix G.. The process involves computing a covariance matrix and using a pre-trained model like GPT-2, leading to dimensions such as $f_{cls}\in\mathbb{R}^{T\in\times 50,257}$, which can be computationally challenging. So, for a model like GPT-2, instead of applying regularization on the output of $f_{cls}$, we apply a linear projection that projects the representation of layer $l$ into a more manageable space, expressed as $X=f_{proj}(f_{l})$, with $f_{proj}\in\mathbb{R}^{T\times 2048}$. This projection layer is inserted over the representation just before the Seq-VCR regularization loss calculation and is trained exclusively end-to-end with the Seq-VCR loss defined below.
Now we can assume that we have $N$ samples in the batch, so $N$ number of $X$, denoted as $\mathbf{X}\in R^{N\times T\times d}$. We can calculate the covariance matrix across the batch dimension, $C\in R^{T\times d\times d}$.
@@
Where for each position in the sequence, covariance matrix can be calculated across the batch dimension when $\mathbf{X}_{:,i}$ is the input for $i^{th}$ position,
$C_{i}(\mathbf{X}_{:,i})=\frac{1}{N-1}\sum_{j=1}^{N}(\mathbf{X}_{j,i}-\bar{\mathbf{X}}_{:,i})(\mathbf{X}_{j,i}-\bar{\mathbf{X}}_{:,i})^{\top},\quad\text{where}\quad\bar{\mathbf{X}}_{:,i}=\frac{1}{N}\sum_{j=1}^{N}\mathbf{X}_{j,i}$ (4)
where $\lambda_{1}$ and $\lambda_{2}$ are the coefficients of the variance and covariance regularization terms. $\eta$ is a small constant (set to 0.001 in our experiments), used for numerical stability. Detailed hyper-parameter choices can be found in Appendix A.
The Variance Term encourages unit variance in each dimension, while the Covariance Term penalizes covariance between different dimensions, promoting de-correlation and diversity in representations.
@@ 3.5 Incorporating Pause Tokens
Increasing the model capacity brings significant accuracy boost for solving $n\times n$ digit multiplication tasks (Qiu et al., 2024). While some previous work increases depth to increase the model capacity, an alternative solution is to use pause tokens that serve as explicit indicators for the model to temporarily pause on intermediate states in sequential tasks before proceeding to the next computation step (Goyal et al., 2023b).
Similarly, we enhance the model’s reasoning capacity by introducing dummy pause tokens, which act as placeholders for intermediate computation steps. This offers notable benefits over Chain-of-Thought (CoT) reasoning tokens, particularly in reducing inference time (as there are often more CoT tokens to decode than pause tokens) and dependency on costly human-supervised data. With pause tokens we can solve tasks like multiplication in a fraction of the time compared to CoT, while performing at a similar accuracy (5 times faster and close to 100% in our experiments. See Appendix E for further details). In all experiments, pause tokens were placed between input and output tokens to emulate CoT reasoning. In particular, the input-output format looked like:
<question> </pause_start> <pause> <pause> </pause_end> <answer>.
We tried with 2, 4, 6, and 8 pause tokens on $4\times 4$ and $5\times 5$ digit multiplication tasks, and we did not find any correlation with task complexity. As such, we used 2 pause tokens in all our experiments. We believe that this may be due to the fact that all pause tokens share the same embedding. Future work will explore the effect of having different embeddings per pause tokens.
@@ 3.6 Training Objective
This encourages the model to not only predict the next token accurately, but also maintain diverse and informative intermediate representations.
@@ 4.1 Data
Finally, we also conduct experiments in a more general setting outside arithmetic tasks, called Dynamic Programming (DP). Dynamic Programming is a framework for solving decision-making problems. The core idea in this setting is to break down a complex task into a series of smaller sub-tasks that can be solved sequentially. We chose a very popular problem in this setting called the Longest Increasing Sub-sequence (LIS) as described in the Introduction to Algorithms book (Cormen et al., 2022). For this task, the goal is to find the length of the longest increasing subsequence of a given integer sequence. We generate datasets with different input sequence lengths ranging from {50, 80, 100} as shown in Feng et al. (2024). Moreover, all input sequences, and answers in LIS are bounded-range integers and can therefore be tokenized (similar to the multi-digit multiplication task). The CoT tokens for this task consist of the dynamic programming array plus the final answer. Figure 3 shows a working example for a sample problem in all the datasets considered.
Table 3 in Appendix D provides an overview of the different datasets i.e., Multiplication, Arithmetic Expression and LIS, along with their respective input, chain-of-thought (CoT), and output token details.
@@ 4.2 Models and Training Configurations
  • Vanilla: Standard training/finetuning without CoT or pause tokens.
  • With CoT: Training/finetuning with explicit CoT supervision.
@@ 4.3.1 Representation Diversity
Seq-VCR significantly enhances representation diversity, as evidenced by the layer-wise distribution of representation entropy across different configurations. In Figure 4, we observe that the configurations that employ our regularization technique, particularly Seq-VCR and Seq-VCR+Pause, exhibit higher entropy values in the intermediate layers compared to the other configurations (Vanilla, and Pause). This increased diversity in the representation entropy suggests that the model is able to capture a wider range of features and maintain distinct representations. The histograms also reveal that while other configurations show peaks at lower entropy values, indicating potential representation collapse and redundancy in the intermediate representations, our method fosters a more uniform distribution across the layers. This uniformity suggests that the model is using a richer feature space, enabling it to better generalize to unseen data, thus improving performance in downstream tasks.
@@ 4.4 Learning Dynamics
Our regularization method, Seq-VCR or Seq-VCR+Pause cause a significant phase transition (improving the next token prediction loss) in the performance of the model across datasets compared to other configurations such as Vanilla, and Pause, which do not induce this phase transition as shown in Figure 5. Specifically, this phase transition marks a distinct shift in the model’s ability to capture and generalize representations during training or fine-tuning.
Without regularization, learning is difficult and training loss saturates. Seq-VCR improves the learning curve. Adding Seq-VCR with $2$ Pause tokens causes a sharp phase transition (Figure 5 (a)) and solves the $5\times 5$ digit multiplication task. Figure 5 (b) shows similar improvements in the arithmetic reasoning task that is learned from scratch.
@@ 4.5 Results on $4\times 4$ and $5\times 5$ digits Multiplication Tasks
We also see from Table 1 that GPT-4 with CoT 5-shot prompts achieves $44.0\%$ for $5\times 5$ digit multiplication, while Vanilla, Seq-VCR, and Pause achieve $0\%$ without CoT. Because, all configurations struggle with predicting middle tokens in the output sequence (shown in Figure 6). This happens because those middle tokens require more computations to calculate successive carry and store them, since these tokens require more computation (as shown in Figure 1), thus making them difficult sub-tasks than predicting the peripheral tokens which require less computations, thus easier to solve. However, Seq-VCR+Pause with $2$ Pause tokens solves the $5\times 5$ digit multiplication task, previously unsolved without CoT tokens. Figure 2(b) shows our regularization Seq-VCR increases intermediate layer entropy, allowing more exploration in the representation space. We hypothesize that along with better representation by Seq-VCR, $2$ Pause tokens increase the computational ability, allowing the model to carry the necessary multiplication values and solve the task. More results by training GPT-2 from scratch are in the Appendix H.
Figure 6: Position-wise accuracy of Multiplication of 5x5 digits with different configuration of GPT-2 Small. We see that models fail on the middle tokens as they require more compute.
@@ 4.6 Results on Arithmetic Expression and Longest Increasing Sub-sequence (LIS) Datasets
Figure 7 illustrates how the accuracy varies among different methods (represented by different bar colors in the plot) and with changing levels of task complexity, such as the number of operators in the Arithmetic Expressions Dataset or the input sequence length in the case of the LIS dataset.
We can observe that as the task complexity is low, i.e. in case of 4 operators in the arithmetic expression and input sequence length of 50 in the LIS dataset, all the methods perform quite well. But, as the number of operators or sequence length increases, the tasks become more complex, which shows the need for models other than the vanilla baselines.
Seq-VCR demonstrates substantial improvement over vanilla models and achieved results comparable to the chain-of-thought (CoT) prompting method as shown in Feng et al. (2024). Although CoT outperforms the other methods, there is an increase in the serial computation due to its step-by-step reasoning approach. On the other hand, Pause tokens and regularization primarily boost parallel computation, enhancing the model’s expressivity for these types of problem without significantly increasing serial computation. The results presented in Figure 7 are based on models with a 5-layer configuration, optimized using the best-performing number of pause tokens. We find that the number of pause tokens that works best for a particular model and layer config vary a bit. For example, the best working Seq-VCR + Pause model uses 7 pause tokens in the case of the 5 operator task in Arithmetic Expression, while 4 pause tokens work the best for the LIS task with 80 as input sequence length. In Figure 8 we show how the different models perform in the most complex task in each dataset. We also clearly see that more depth of the model helps every method in solving the task better. Another conclusion we can draw from Figures 7 and 8 is that although the combination of Seq-VCR with Pause Tokens works best (almost close to the CoT method), Seq-VCR achieves a similar test accuracy without the use of Pause Tokens in both datasets. All results in Figures 7 and 8 are over 3 seed runs. Detailed ablation studies, exploring variations in the number of pause tokens, layers, and task complexities, can be found in the Appendices  B and  C.